OpenStudy (anonymous):

I need help i really don't understand this thing i will fan and give a medal for who ever helps

OpenStudy (anonymous):

i have to fine the x for a rhombus hold on ill upload a picture

OpenStudy (anonymous):

ill show you what i did and where i got lost

OpenStudy (anonymous):

ok i suck at this but ill try to help

OpenStudy (anonymous):

okay this is what i did (4x^2-24)=(3x^2-5x) x^2-24=-5x x^2-5x-24=0 i did all this but ill show you what my study plan show afterwards

OpenStudy (anonymous):

i dont know what to do after ive done this

OpenStudy (anonymous):

@mz_samosino are you smarter then me?

OpenStudy (anonymous):

x^2-5x-24=0 are you looking for factors?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x-3)(x+8) x=-8 , 3

OpenStudy (anonymous):

how did you get that? did you divide

OpenStudy (anonymous):

okay nevermind i see it thanks guys for helping me :)

OpenStudy (anonymous):

@kag u suprise me with your question hahah!! :D